# Inclusions of real rank zero

### v2

@@ Abstract.
The first named author was supported by the IRFD grants 1054-00094B and 1026-00371B. The second named author was supported by the EPSRC grant EP/R513295/1.
@@ Introduction
The real rank of $\mathrm{C}^{*}$-algebras was introduced by Brown and Pedersen in [10] as a non-commutative analogue of topological dimension. Particularly, the notion of real rank zero distinguished itself due to the fact that many interesting classes of $\mathrm{C}^{*}$-algebras were shown to have this property. To mention a few, simple purely infinite $\mathrm{C}^{*}$-algebras ([54]), irrational rotation algebras ([14]), AF-algebras ([10]), and von Neumann algebras ([10]) have real rank zero. Subsequently, real rank zero started to appear in the context of the classification programme of simple nuclear $\mathrm{C}^{*}$-algebras. Most notably in [20], real rank zero facilitated access to classification by $K$-theory for large classes of $\mathrm{C}^{*}$-algebras.
@@
On the other hand, the study of inclusions in operator algebras is ubiquitous in the last four decades. In the case of von Neumann algebras we can mention Jones’ celebrated theory of subfactors ([28]), or the Cartan subalgebras of Feldman–Moore ([21]). On the $\mathrm{C}^{*}$-algebraic side, inclusions appear in many concrete realisations, such as the AF core of a Cuntz-Krieger algebra ([16]) or Cartan pairs induced by an étale groupoid (cf. [40]), and more abstractly in Watatani’s index theory ([50]) or Rørdam’s $\mathrm{C}^{*}$-irreducible inclusions ([44]). Inclusions also appear implicitly in the Elliott classification programme, as one often classifies inclusions of simple nuclear $\mathrm{C}^{*}$-algebras up to approximate unitary equivalence ([22, 46, 23, 24, 26, 11]).
@@
Furthermore, there are natural inclusions arising from dynamics which have real rank zero. The canonical inclusion $\mathrm{C}^{*}_{r}(G)\subseteq\ell^{\infty}(G)\rtimes_{r}G$ has real rank zero (Theorem 5.7) for any countable discrete exact group $G$. This example is particularly intriguing since real rank zero seems to be quite a rare phenomenon both for group $\mathrm{C}^{*}$-algebras and for uniform Roe algebras. [^1] We invite the reader to see the discussion in [35, Section 3]. For example, the uniform Roe algebra of $SL(n,\mathbb{Z})$ does not have real rank zero if $n\geq 3$.
@@
Let $A\subseteq B$ be an inclusion of separable exact $\mathrm{C}^{*}$-algebras such that $1_{B}\in A$ and suppose that the inclusion $\iota\colon A\hookrightarrow B$ is full, nuclear and $\mathcal{O}_{\infty}$-stable. Then $\iota$ is approximately unitarily equivalent to a $*{}^{*}$-homomorphism which factors through a $\mathrm{C}^{*}$-algebra with real rank zero.
This paper is organised as follows. In Section 1 we introduce real rank zero for inclusions and $*{}^{*}$-homomorphisms. After showing a series of equivalent characterisations for this property, we offer a complete description of the commutative case in Section 2. Then, Section 3 collects a number of permanence properties. In Section 4 we examine inclusions of real rank zero at the level of their invariant. In particular, we prove that under some regularity assumptions, real rank zero can be characterised by the relative pairing of $K$-theory and traces. Then, Section 5 focuses on infinite inclusions and provides some non-trivial examples of inclusions of real rank zero.
@@ Acknowledgements
The first named author would like to thank Matt Kennedy for suggesting applications to Furstenberg boundaries. The second named author would like to thank Stuart White for his supervision on this project, Rufus Willett and Julian Kranz for helpful conversations on boundaries of hyperbolic groups, and Alistair Miller for asking a question that helped extending the result in Theorem 5.7. This article will form part of the DPhil thesis of the second named author.
@@ 1. Equivalence of definitions
A natural source of inclusions arises from $*{}^{*}$-homomorphisms by considering the inclusion of the image into the codomain. Then, if this inclusion has real rank zero, we will say that the $*{}^{*}$-homomorphism has real rank zero.
@@
Let $\theta:A\to B$ be a $*{}^{*}$-homomorphism between $\mathrm{C}^{*}$-algebras. Then we say that $\theta$ has real rank zero if the inclusion $\theta(A)\subseteq B$ has real rank zero.
In the spirit of [10, Theorem 2.6], let us show that there are many equivalent characterisations of when an inclusion has real rank zero. Since some of the properties require the existence of a unit, let us first prove the following theorem in the unital case.
@@
We are going to show that (iii) and (iv) are equivalent, and then that (i)$\implies$(iv)$\implies$(ii)$\implies$ (i).
Assume condition (iii). We will show that condition (iv) holds. Let $x,y\in A$ be positive and mutually orthogonal, $\epsilon>0$, and $\epsilon_{0}>0$ such that $\epsilon_{0}+\sqrt{(4+\epsilon_{0})\epsilon_{0})}<\epsilon$. If $x=0$, we take $p=0$ and if $y=0$, we take $p=1$. Therefore, we can assume that $x$ and $y$ are nonzero. Rescalling if necessary, suppose that $\|x\|=\|y\|=1.$ If we let $a=x-y$ which is self-adjoint, then by (iii), there exists $b\in B$ invertible and self-adjoint such that $\|a-b\|\leq\epsilon_{0}$. Let $p$ be the spectral projection of $b$ supported on $(0,\infty]$. Then, we note that $pb_{+}=b_{+}$ and $(1-p)b_{-}=b_{-}$. [^3] Note that $b_{+}=\max(b,0)$ and $b_{-}=\max(-b,0)$. In particular, $b=b_{+}-b_{-}$. We claim that $p$ satisfies the required conditions in (iv). Since $xy=0$, it follows that $a_{+}=(x-y)_{+}=x$. As $pb_{+}=b_{+}$, by an application of the triangle inequality, we have that
@@
By [10, Lemma $2.2$], we get that $2\|a_{+}-b_{+}\|\leq\delta+\epsilon_{0}$, where $\delta^{2}=(\|a\|+\|b\|)\epsilon_{0}$. Since $\|a\|\leq 2$, we have that $\|b\|\leq 2+\epsilon_{0}$ and hence $\delta^{2}\leq(4+\epsilon_{0})\epsilon_{0}$. Therefore,
@@
Since $\|pa_{-}p\|\leq\epsilon$, we have $pa_{-}p\leq\epsilon p$, so $pap\geq p(a_{+}-\epsilon)p\geq-\epsilon p$. Therefore, we get that $pbp\geq\epsilon p$, which shows that $pbp$ is invertible in $pBp$. Similarly, $(1-p)a(1-p)\leq\epsilon(1-p)$, so $(1-p)b(1-p)\leq-\epsilon(1-p)$, which gives that $(1-p)b(1-p)$ is invertible in $(1-p)B(1-p)$. Thus, $b$ is invertible in $B$, by an application of [10, Lemma 2.3] with $x=b$ and $c=d=0$. Finally, using (1.1), a standard application of the triangle inequality shows that $\|a-b\|\leq 8\epsilon$.
We then move to the implication (i)$\implies$(iv). Let $x,y\in A$ positive orthogonal and $\epsilon>0$. Since $A\subseteq B$ has real rank zero, there exists a projection $p\in\overline{xBx}$ such that $\|px-x\|\leq\epsilon$. Since $xy=0$, we also get that $py=0$, so (iv) holds.
@@
As the spectrum of $p_{j}a$ in $D_{j}:=C^{*}((a-j/n)_{+},p_{j})$ is contained in $[j/n,1]$ and $p_{j}$ commutes with $a$, we get that $p_{j}=f_{j}(p_{j}a)$. Moreover, since the map $\phi:\mathrm{C}^{*}(a)\to D_{j}$ given by $x\mapsto p_{j}x$ is a $*{}^{*}$-homomorphism, we also get that $p_{j}f_{j}(a)=f_{j}(p_{j}a)=f_{j}(ap_{j})=f_{j}(a)p_{j}$.
@@
Since $(q_{j})_{j=0}^{n-1}$ is a sequence of mutually orthogonal projections that sum up to $1_{B_{\infty}}$, the map $\phi:\mathbb{C}^{n}\to B_{\infty}$ given by $(\lambda_{0},\ldots,\lambda_{n-1})\mapsto\sum\limits_{j=0}^{n-1}\lambda_{j}q_{j}$ is a unital $*{}^{*}$-homomorphism. Let $c_{j}(B)=\{(b_{i})_{i\geq 1}\in\ell^{\infty}(B):b_{i}=0\ \forall\ i\geq j\}$ and note that $(c_{j}(B))_{j\geq 1}$ is an increasing sequence of ideals of $\ell^{\infty}(B)$ generating $c_{0}(B)$. Then, since $\mathbb{C}^{n}$ is semiprojective ([36, Lemma 14.1.5] and [36, Theorem 14.2.1]), it follows that $\phi$ lifts to a unital $*{}^{*}$-homomorphism $\tilde{\phi}:\mathbb{C}^{n}\to\ell^{\infty}(B)/c_{j}(B)\subseteq\ell^{\infty}(B)$ for some $j\geq 1$. Say that $\tilde{\phi}$ is given by a sequence of $*{}^{*}$-homomorphisms $(\phi_{k})_{k\geq 1}$. Then $x_{k}=\phi_{k}(0,1/n,\ldots,(n-1)/n)$ is a self-adjoint element with finite spectrum and (1.6) gives that $\limsup\limits_{k\to\infty}\|a-x_{k}\|\leq 2/n$. Thus, for $k$ large, we get that $\|a-x_{k}\|\leq 3/n<\epsilon$, so condition (ii) holds as claimed.
Suppose now that (ii) holds and we want to obtain (i). We remark for the reader’s convenience that this part of the proof does not use the assumption that the inclusion is unital. Let $a\in A_{+}$ be nonzero, and $E$ be the hereditary $\mathrm{C}^{*}$-subalgebra of $B$ generated by $a$. Rescalling if necessary, we can assume that $\|a\|=1$. Given $10<\epsilon<$ 1, it suffices to find a projection $p\in E$ such that $\|a-pa\|\leq 2\epsilon$. Choose $\delta>0$ such that $9\delta<\epsilon-\epsilon^{2}$ and $n\in\mathbb{N}$ such that $1-\delta^{2/n}\leq\delta$.
@@
If we denote by $q$ the spectral projection of $b$ corresponding to the interval $[\delta,1]$, then $\|b-qb\|\leq\delta$. Moreover, $q\in B$ as $b$ has finite spectrum. Recall that $1-\delta^{2/n}\leq\delta$. Since the spectrum of $bq$ in $qBq$ is contained in $[\delta,1]$, it follows that
@@
and by our choice of $\delta$, the spectrum of $z$ is contained in $[0,\epsilon]\cup[1-\epsilon,1]$. If we consider the spectral projection $p$ of $z$ corresponding to the interval $[1/2,1]$, then $p\in E$ as $z\in E$, and $\|z-p\|\leq\epsilon$. Hence, $\|p-q\|\leq\epsilon+3\delta$ and thus
@@
Following the remark in the proof of Theorem 1.6, the fact that (ii) implies (i) is obtained in the proof of (ii)$\implies$(i) in Theorem 1.6. Then, it only remains to prove that (iv) implies (ii). Let $a\in A$ be self-adjoint, $\epsilon>0$ and suppose that (iv) holds. Then, applying Theorem 1.6 to the inclusion $A^{\dagger}\subseteq B^{\dagger}$, there exists $b\in B^{\dagger}$ self-adjoint with finite spectrum such that $\|a-b\|\leq\epsilon/2$. Then $b=b_{0}+\lambda 1$, so $b_{0}\in B$ is self-adjoint, has finite spectrum, and $\|a-b_{0}\|\leq\epsilon$. Hence, the conclusion follows. ∎
@@
Let $A,B$, and $C$ be $\mathrm{C}^{*}$-algebras. Then the set of $*{}^{*}$-homomorphisms from $A$ to $B$ with real rank zero is point-norm closed. Moreover, if $\phi:C\to B$ and $\psi:A\to C$ are $*{}^{*}$-homomorphisms, at least one with real rank zero, then $\phi\circ\psi$ has real rank zero.
@@
Let $\theta_{n}:A\to B$ be a sequence of $*{}^{*}$-homomorphisms with real rank zero converging pointwise to $\theta:A\to B$. Then $\theta$ has real rank zero by using condition (ii) in Theorem 1.7 for the sequence of inclusions $\theta_{n}(A)\subseteq B$. Suppose now $\phi:C\to B$ and $\psi:A\to C$ are $*{}^{*}$-homomorphisms and $\psi$ has real rank zero. Let $a\in A$ be self-adjoint and $\epsilon>0$. Then there exists $c\in C$ self-adjoint with finite spectrum such that $\|\psi(a)-c\|\leq\epsilon$. Therefore, $\phi(c)$ is a self-adjoint element with finite spectrum and $\|\phi(\psi(a))-\phi(c)\|\leq\epsilon$. Hence $\phi\circ\psi$ has real rank zero. The case when $\phi$ has real rank zero follows similarly. ∎
@@
Let $C$ be a $\mathrm{C}^{*}$-algebra of real rank zero, $\psi:A\to C$ and $\phi:C\to B$ be $*{}^{*}$-homomorphisms, and $(u_{\lambda})_{\lambda}$ be a net of unitaries in the multiplier algebra of $B$ such that $\operatorname{Ad}(u_{\lambda})\circ\phi\circ\psi$ converges pointwise to the inclusion $A\subseteq B$.
Since $C$ has real rank zero, combining Lemma 1.9 and Lemma 1.10 shows that the net of $*{}^{*}$-homomorphisms $\operatorname{Ad}(u_{\lambda})\circ\phi\circ\psi$ have real rank zero. Then, by the first part of the Lemma 1.10, we get that the inclusion $A\subseteq B$ has real rank zero. ∎

### v3

@@ Introduction
The real rank of $\mathrm{C}^{*}$-algebras was introduced by Brown and Pedersen in [11] as a non-commutative analogue of topological dimension. Particularly, the notion of real rank zero distinguished itself due to the fact that many interesting classes of $\mathrm{C}^{*}$-algebras were shown to have this property. To mention a few, simple purely infinite $\mathrm{C}^{*}$-algebras ([56]), irrational rotation algebras ([6, 21]), AF-algebras ([11]), and von Neumann algebras ([11]) have real rank zero. Subsequently, real rank zero started to appear in the context of the classification programme of simple nuclear $\mathrm{C}^{*}$-algebras. Most notably in [20], real rank zero facilitated access to classification by $K$-theory for large classes of $\mathrm{C}^{*}$-algebras.
@@
On the other hand, the study of inclusions in operator algebras is ubiquitous in the last four decades. In the case of von Neumann algebras we can mention Jones’ celebrated theory of subfactors ([30]), or the Cartan subalgebras of Feldman–Moore ([22, 23]). On the $\mathrm{C}^{*}$-algebraic side, inclusions appear in many concrete realisations, such as the AF core of a Cuntz-Krieger algebra ([16]) or Cartan pairs induced by an étale groupoid (cf. [42]), and more abstractly in Watatani’s index theory ([52]) or Rørdam’s $\mathrm{C}^{*}$-irreducible inclusions ([46]). Inclusions also appear implicitly in the Elliott classification programme, as one often classifies inclusions of simple nuclear $\mathrm{C}^{*}$-algebras up to approximate unitary equivalence ([24, 48, 25, 26, 28, 12]).
@@
Furthermore, there are natural inclusions arising from dynamics which have real rank zero. The canonical inclusion $\mathrm{C}^{*}_{r}(G)\subseteq\ell^{\infty}(G)\rtimes_{r}G$ has real rank zero (Theorem 5.6) for any countable discrete exact group $G$. This example is particularly intriguing since real rank zero seems to be quite a rare phenomenon both for group $\mathrm{C}^{*}$-algebras and for uniform Roe algebras. [^1] We invite the reader to see the discussion in [37, Section 3]. For example, the uniform Roe algebra of $SL(n,\mathbb{Z})$ does not have real rank zero if $n\geq 3$.
@@
Let $A\subseteq B$ be an inclusion of separable exact $\mathrm{C}^{*}$-algebras such that $1_{B}\in A$ and suppose that the inclusion $\iota\colon A\hookrightarrow B$ is full, nuclear and $\mathcal{O}_{\infty}$-stable. Then $\iota$ is approximately unitarily equivalent to a -homomorphism which factors through a $\mathrm{C}^{*}$-algebra with real rank zero.
This paper is organised as follows. In Section 1 we introduce real rank zero for inclusions and -homomorphisms. After showing a series of equivalent characterisations for this property, we offer a complete description of the commutative case in Section 2. Then, Section 3 collects a number of permanence properties. In Section 4 we examine inclusions of real rank zero at the level of their invariant. In particular, we prove that under some regularity assumptions, real rank zero can be characterised by the relative pairing of $K$-theory and traces. Then, Section 5 focuses on infinite inclusions and provides some non-trivial examples of inclusions of real rank zero.
@@ Acknowledgements
The first-named author was supported by the IRFD grants 1054-00094B and 1026-00371B. The second-named author was supported by the EPSRC grant EP/R513295/1. The first-named author would like to thank Matt Kennedy for suggesting applications to Furstenberg boundaries. The second-named author would like to thank Stuart White for his supervision on this project, Rufus Willett and Julian Kranz for helpful conversations on boundaries of hyperbolic groups, and Alistair Miller for asking a question that helped extending the result in Theorem 5.6. This article forms part of the DPhil thesis of the second-named author.
@@ 1. Equivalence of definitions
A natural source of inclusions arises from -homomorphisms by considering the inclusion of the image into the codomain. Then, if this inclusion has real rank zero, we will say that the -homomorphism has real rank zero.
@@
Let $\theta:A\to B$ be a -homomorphism between $\mathrm{C}^{*}$-algebras. Then we say that $\theta$ has real rank zero if the inclusion $\theta(A)\subseteq B$ has real rank zero.
In the spirit of [11, Theorem 2.6], let us show that there are many equivalent characterisations of when an inclusion has real rank zero. Since some of the properties require the existence of a unit, let us first prove the following theorem in the unital case.
@@
[…]
Assume condition (iii). We will show that condition (iv) holds. Let $x,y\in A$ be positive and mutually orthogonal, $\epsilon>0$, and $\epsilon_{0}>0$ such that $\epsilon_{0}+\sqrt{(4+\epsilon_{0})\epsilon_{0})}<\epsilon$. If $x=0$, we take $p=0$ and if $y=0$, we take $p=1$. Therefore, we can assume that $x$ and $y$ are nonzero. Rescaling if necessary, suppose that $\|x\|=\|y\|=1.$ If we let a=xya=x-y which is self-adjoint, then by (iii), there exists $b\in B$ invertible and self-adjoint such that $\|a-b\|\leq\epsilon_{0}$. Let $p$ be the spectral projection of $b$ supported on $(0,\infty]$. Then, we note that $pb_{+}=b_{+}$ and $(1-p)b_{-}=b_{-}$. [^3] Note that $b_{+}=\max(b,0)$ and $b_{-}=\max(-b,0)$. In particular, $b=b_{+}-b_{-}$. We claim that $p$ satisfies the required conditions in (iv). Since $xy=0$, it follows that $a_{+}=(x-y)_{+}=x$. As $pb_{+}=b_{+}$, by an application of the triangle inequality, we have that
@@
By [11, Lemma $2.2$], we get that $2\|a_{+}-b_{+}\|\leq\delta+\epsilon_{0}$, where $\delta^{2}=(\|a\|+\|b\|)\epsilon_{0}$. Since $\|a\|\leq 2$, we have that $\|b\|\leq 2+\epsilon_{0}$ and hence $\delta^{2}\leq(4+\epsilon_{0})\epsilon_{0}$. Therefore,
@@
Since $\|pa_{-}p\|\leq\epsilon$, we have $pa_{-}p\leq\epsilon p$, so $pap\geq p(a_{+}-\epsilon)p\geq-\epsilon p$. Therefore, we get that $pbp\geq\epsilon p$, which shows that pbppbp is invertible in $pBp$. Similarly, $(1-p)a(1-p)\leq\epsilon(1-p)$, so $(1-p)b(1-p)\leq-\epsilon(1-p)$, which gives that (1p)b(1p)(1-p)b(1-p) is invertible in (1p)B(1p)(1-p)B(1-p). Thus, $b$ is invertible in $B$, by an application of [11, Lemma 2.3] with $x=b$ and $c=d=0$. Finally, using (1.1), a standard application of the triangle inequality shows that $\|a-b\|\leq 8\epsilon$.
[…]
@@
As the spectrum of $p_{j}a$ in $D_{j}:=C^{*}((a-j/n)_{+},p_{j})$ is contained in $[j/n,1]$ and $p_{j}$ commutes with $a$, we get that $p_{j}=f_{j}(p_{j}a)$. Moreover, since the map $\phi:\mathrm{C}^{*}(a)\to D_{j}$ given by $x\mapsto p_{j}x$ is a -homomorphism, we also get that $p_{j}f_{j}(a)=f_{j}(p_{j}a)=f_{j}(ap_{j})=f_{j}(a)p_{j}$.
@@
Since $(q_{j})_{j=0}^{n-1}$ is a sequence of mutually orthogonal projections that sum up to $1_{B_{\infty}}$, the map $\phi:\mathbb{C}^{n}\to B_{\infty}$ given by $(\lambda_{0},\ldots,\lambda_{n-1})\mapsto\sum\limits_{j=0}^{n-1}\lambda_{j}q_{j}$ is a unital -homomorphism. Let $c_{j}(B)=\{(b_{i})_{i\geq 1}\in\ell^{\infty}(B):b_{i}=0\ \forall\ i\geq j\}$ and note that $(c_{j}(B))_{j\geq 1}$ is an increasing sequence of ideals of $\ell^{\infty}(B)$ generating $c_{0}(B)$. Then, since $\mathbb{C}^{n}$ is semiprojective ([38, Lemma 14.1.5] and [38, Theorem 14.2.1]), it follows that $\phi$ lifts to a unital -homomorphism $\tilde{\phi}:\mathbb{C}^{n}\to\ell^{\infty}(B)/c_{j}(B)\subseteq\ell^{\infty}(B)$ for some $j\geq 1$. Say that $\tilde{\phi}$ is given by a sequence of -homomorphisms $(\phi_{k})_{k\geq 1}$. Then $x_{k}=\phi_{k}(0,1/n,\ldots,(n-1)/n)$ is a self-adjoint element with finite spectrum and (1.6) gives that $\limsup\limits_{k\to\infty}\|a-x_{k}\|\leq 2/n$. Thus, for $k$ large, we get that $\|a-x_{k}\|\leq 3/n<\epsilon$, so condition (ii) holds as claimed.
Suppose now that (ii) holds and we want to obtain (i). We remark for the reader’s convenience that this part of the proof does not use the assumption that the inclusion is unital. Let $a\in A_{+}$ be nonzero, and $E$ be the hereditary $\mathrm{C}^{*}$-subalgebra of $B$ generated by $a$. Rescaling if necessary, we can assume that $\|a\|=1$. Given $0<\epsilon<1$, it suffices to find a projection $p\in E$ such that $\|a-pa\|\leq 2\epsilon$. Choose $\delta>0$ such that $9\delta<\epsilon-\epsilon^{2}$ and $n\in\mathbb{N}$ such that $1-\delta^{2/n}\leq\delta$.
@@
If we denote by $q$ the spectral projection of $b$ corresponding to the interval $\delta,1]$, then $\|b-qb\|\leq\delta$. Moreover, $q\in B$ as $b$ has finite spectrum. Recall that $1-\delta^{2/n}\leq\delta$. Since the spectrum of $bq$ in $qBq$ is contained in $\delta,1]$, it follows that
@@
and by our choice of $\delta$, the spectrum of $z$ is contained in $0,\epsilon]\cup[1-\epsilon,1]$. If we consider the spectral projection $p$ of $z$ corresponding to the interval $[1/2,1]$, then $p\in E$ as $z\in E$, and $\|z-p\|\leq\epsilon$. Hence, $\|p-q\|\leq\epsilon+3\delta$ and thus
@@
Following the remark in the proof of Theorem 1.6, the fact that (ii) implies (i) is obtained in the proof of $\implies(i) i$n Theorem 1.6. Then, it only remains to prove that (iv) implies (ii). Let $a\in A$ be self-adjoint, $\epsilon>0$ and suppose that (iv) holds. Then, applying Theorem 1.6 to the inclusion $A^{\dagger}\subseteq B^{\dagger}$, there exists $b\in B^{\dagger}$ self-adjoint with finite spectrum such that $\|a-b\|\leq\epsilon/2$. Then $b=b_{0}+\lambda 1$, so $b_{0}\in B$ is self-adjoint, has finite spectrum, and $\|a-b_{0}\|\leq\epsilon$. Hence, the conclusion follows. ∎
@@
Let $A,B$, and $C$ be $\mathrm{C}^{*}$-algebras. Then the set of -homomorphisms from $A$ to $B$ with real rank zero is point-norm closed. Moreover, if $\phi:C\to B$ and $\psi:A\to C$ are -homomorphisms, at least one with real rank zero, then $\phi\circ\psi$ has real rank zero.
@@
Let $\theta_{n}:A\to B$ be a sequence of -homomorphisms with real rank zero converging pointwise to $\theta:A\to B$. Then $\theta$ has real rank zero by using condition (ii) in Theorem 1.7 for the sequence of inclusions $\theta_{n}(A)\subseteq B$. Suppose now $\phi:C\to B$ and $\psi:A\to C$ are -homomorphisms and $\psi$ has real rank zero. Let $a\in A$ be self-adjoint and $\epsilon>0$. Then there exists $c\in C$ self-adjoint with finite spectrum such that $\|\psi(a)-c\|\leq\epsilon$. Therefore, $\phi(c)$ is a self-adjoint element with finite spectrum and $\|\phi(\psi(a))-\phi(c)\|\leq\epsilon$. Hence $\phi\circ\psi$ has real rank zero. The case when $\phi$ has real rank zero follows similarly. ∎
@@
Let $C$ be a $\mathrm{C}^{*}$-algebra of real rank zero, $\psi:A\to C$ and $\phi:C\to B$ be -homomorphisms, and $(u_{\lambda})_{\lambda}$ be a net of unitaries in the multiplier algebra of $B$ such that $\operatorname{Ad}(u_{\lambda})\circ\phi\circ\psi$ converges pointwise to the inclusion $A\subseteq B$.
Since $C$ has real rank zero, combining Lemma 1.9 and Lemma 1.10 shows that the net of -homomorphisms $\operatorname{Ad}(u_{\lambda})\circ\phi\circ\psi$ have real rank zero. Then, by the first part of the Lemma 1.10, we get that the inclusion $A\subseteq B$ has real rank zero. ∎